You win Rs 10,000 in a horse race. You have a choice between spending the money now and putting it in a 5% FD. What is the opportunity cost of spending the Rs 10,000 now?
1
Expert's answer
2021-10-10T16:34:54-0400

Solution:

Opportunity cost refers to the value of the next best-foregone alternative. It is the benefit that would have been obtained if a different option had been chosen.

When you invest in a Fixed Deposit account, the value will be as follows:

Return = 10,000 + (10,000 x 5% x 1) =10,000 + 500 = Rs 10,500


Opportunity cost = Rs 10,500 since it is the value of the next best-foregone option.


The opportunity cost of spending the Rs 10,000 now is Rs 10,500.